# Classifying destructive quantum interference in molecular junctions: Towards molecular quantum rulers

### v1

@@ II Transport description
is the spectral representation of the zeroth-order retarded Green’s function. The sum runs over all molecular orbitals $n$ with energy $\epsilon_{n}$, which might be considered to be renormalized by the presence of the electrodes. The expansion coefficients of the $n$-th orbital at the anchoring molecular sites ‘l’ and ‘r’ are denoted by $c_{\text{l},n}$ and $c_{\text{r},n}$, respectively, and the star indicates the complex conjugate. The product of the orbital expansion coefficients is abbreviated as the residue $\rho_{n}=c_{\text{l},n}c^{*}_{\text{r},n}$. The imaginary part $\eta$ can be viewed to arise from the finite lifetime acquired by the molecule-electrode coupling. For numerical purposes, we choose $\eta=10^{-4}\Delta$ to avoid divergences of the Green’s function $G_{\text{lr}}^{\text{r}}$, with $\Delta$ being basically the size of the HOMO-LUMO gap. For its precise definition within the four-level model, see Eqs. (4) and (5). Within the analytical treatment, $\eta$ will be set to zero, and to indicate this, we will then drop the superscript ‘r’ and denote the propagator as $G_{\text{lr}}$. As far as the real part of the self-energies is concerned, we assume that the energies $\epsilon_{n}$ refer to the dressed levels of a molecule attached to two leads. We refer the reader to Sec. A for further explanations of approximations involved.
@@ IV Reduction to the four-level model
In the following, we will focus on the case of four orbitals in the sum of Eq. (3) and explore the emerging rich structure of zeros of the transmission inside the HOMO-LUMO gap. The choice of four levels is motivated by the architecture of mechanoelectrically sensitive molecules.Nozaki, Lücke, and Schmidt (2017); Stefani et al. (2018); Reznikova et al. (2021); Schosser et al. (2022); Hsu et al. (2022); van der Poel et al. (2024) They typically consist of two identical conjugated molecular decks that can be mechanically shifted against each other. The weak tunnel-splitting of the HOMO and LUMO of each deck, arising from the coupling of $\pi$ electrons, leads to four relevant molecular states. Their energetic order can be varied, since the sign of the tunnel-splitting of HOMOs and LUMOs depends on the displacement of the molecular decks.Nozaki, Lücke, and Schmidt (2017); Stefani et al. (2018) Further energy levels can be neglected, provided that the applied voltage and the temperature are sufficiently small. In the following, we will collectively refer to HOMO-1 and HOMO as HOMOs and correspondingly for LUMOs. We will also assume that the residues $\rho_{n}$ of all levels are nonzero, as otherwise their number would effectively be less than four.
@@
for the two lowest unoccupied orbitals. We choose $a_{\text{L}}+a_{\text{H}}<1$ to ensure that the HOMO-LUMO gap $\Delta-a_{\text{H}}-a_{\text{L}}$ is always positive. The zero of energy always lies in the middle of the HOMO-LUMO gap. For convenience, we vary the energies periodically as a function of the control parameter as
@@ VI Nonbinary four-level model
So far, we have taken the residues to be independent of the control parameter $x$. In general, this will not be the case. Without an appropriate model for a concrete situation, we limit ourselves to present one example, which exhibits an interesting structure in the HOMO-LUMO gap. In Fig. 9, we have chosen residues to oscillate as a function of the control parameter. The choice of parameters is given in the figure caption. In this specific example, we find a transmission valley with a loop structure as a function $E$ and $x$ near $x=0$. Furthermore, a transmission valley of type III connects degeneracies in the HOMOs and the LUMOs. However, it exhibits a non-monotonic behavior as a function of $x$ and does not connect the two closest degeneracies.
@@ VII Conclusions
Complex destructive quantum interference structures could be used to build precise molecular quantum rulers to measure both energetic and control parameter separations, e.g. changes in electrode displacement, using the molecule inside a molecular junction as a quantum sensor. In this way, much might be learned about fluctuations that single-molecule junctions are subject to. van der Poel et al. (2024)
It is presently not clear in all the cases, how a molecule would need to look like that shows the desired transmission properties. We leave this as a challenge to talented synthetic chemists and experimentalists.

### v2

@@ II Transport description
is the spectral representation of the zeroth-order retarded Green’s function. The sum runs over all molecular orbitals $n$ with energy $\epsilon_{n}$, which might be considered to be renormalized by the presence of the electrodes. The expansion coefficients of the $n$-th orbital at the anchoring molecular sites ‘l’ and ‘r’ are denoted by $c_{\text{l},n}$ and $c_{\text{r},n}$, respectively, and the star indicates the complex conjugate. The product of the orbital expansion coefficients is abbreviated as the residue $\rho_{n}=c_{\text{l},n}c^{*}_{\text{r},n}$. The imaginary part $\eta$ can be viewed to arise from the finite lifetime acquired by the molecule-electrode coupling. For numerical purposes, we choose $\eta=10^{-4}\Delta$ to avoid divergences of the Green’s function $G_{\text{lr}}^{\text{r}}$, with $\Delta$ being basically the size of the HOMO-LUMO gap. For its precise definition within the four-level model, see Eqs. (4) and (5). Within the analytical treatment, $\eta$ will be set to zero, and to indicate this, we will then drop the superscript ‘r’ and denote the propagator as $G_{\text{lr}}$. While a finite value of $\eta$ will lead to a broadening, the features in the transmission will be robust as long as $\eta\ll\Delta$. As far as the real part of the self-energies is concerned, we assume that the energies $\epsilon_{n}$ refer to the dressed levels of a molecule attached to two leads. We refer the reader to Sec. A for further explanations of approximations involved.
@@ IV Reduction to the four-level model
In the following, we will focus on the case of four orbitals in the sum of Eq. (3) and explore the emerging rich structure of zeros of the transmission inside the HOMO-LUMO gap. The choice of four levels is motivated by the architecture of mechanoelectrically sensitive molecules.Nozaki, Lücke, and Schmidt (2017); Stefani et al. (2018); Reznikova et al. (2021); Schosser et al. (2022); Hsu et al. (2022); van der Poel et al. (2024) They typically consist of two identical conjugated molecular decks that can be mechanically shifted against each other. The weak tunnel-splitting of the HOMO and LUMO of each deck, arising from the coupling of $\pi$ electrons, leads to four relevant molecular states. Their energetic order can be varied, since the sign of the tunnel-splitting of HOMOs and LUMOs depends on the displacement of the molecular decks.Nozaki, Lücke, and Schmidt (2017); Stefani et al. (2018) Further energy levels can be neglected, provided that the applied voltage and the temperature are sufficiently small. Below, we will collectively refer to HOMO-1 and HOMO as HOMOs and correspondingly for LUMOs. We will also assume that the residues $\rho_{n}$ of all levels are nonzero, as otherwise their number would effectively be less than four.
@@
for the two lowest unoccupied orbitals. We choose $a_{\text{L}}+a_{\text{H}}<1$ to ensure that the HOMO-LUMO gap $\Delta(1-a_{\text{H}}-a_{\text{L}})$ is always positive. The zero of energy always lies in the middle of the HOMO-LUMO gap. For convenience, we vary the energies periodically as a function of the control parameter as
@@ VI Nonbinary four-level model
So far, we have taken the residues to be independent of the control parameter $x$. In general, this will not be the case. Without an appropriate model for a concrete situation, we limit ourselves to present one example, which exhibits an interesting structure in the HOMO-LUMO gap. In Fig. 9, we have chosen residues to oscillate as a function of the control parameter. The choice of parameters is given in the figure caption. In this specific example, we find a transmission valley with a loop structure as a function of $E$ and $x$ near $x=0$. Furthermore, a transmission valley of type III connects degeneracies in the HOMOs and the LUMOs. However, it exhibits a non-monotonic behavior as a function of $x$ and does not connect the two closest degeneracies.
@@ VII Conclusions
Complex destructive quantum interference structures could be used to build precise molecular quantum rulers to electrically measure both energetic and control parameter separations, e.g. changes in electrode displacement, using the molecule inside a molecular junction as a quantum sensor. In this way, much might be learned about fluctuations that single-molecule junctions are subject to. van der Poel et al. (2024)
It is presently not clear in all the cases, how a molecule would need to look like that shows the desired transmission properties. Destructive quantum interferences that cross the entire HOMO-LUMO gap (type III in Fig. 3) have been reported for $\pi$-stacked molecules early on.Nozaki, Lücke, and Schmidt (2017); Stefani et al. (2018) In simulations of anthracene double-deckers with various anchor group attachment points using extended Hückel parameters, we find between zero and two destructive quantum interferences inside the HOMO-LUMO gap for selected electrode separations, and some destructive interference valleys are loop-like (type II in Fig. 3). The choice of appropriate intermolecular bridgesSchosser et al. (2022) or the attachment points of anchoring groupsHsu et al. (2022) may offer ways to steer the sliding motion of the $\pi$-electron systems of two molecular decks. An unfolding, as in helicene spirals, may additionally allow for changing orbital orientation upon stretching, offering possibilities to vary residue sizes. We hope that the predicted quantum interference types will be observed in the future. The finding of suitable molecular structures represents an exciting challenge for talented synthetic chemists and experimentalists.